# Efficient and Accurate Spatial Mixing of Machine Learned Interatomic Potentials for Materials Science

### v1

@@ Abstract
Machine-learned interatomic potentials offer near first-principles accuracy but are computationally expensive, limiting their application in large-scale molecular dynamics simulations. Inspired by quantum mechanics/molecular mechanics methods, we present ML-MIX, an efficient and flexible LAMMPS package for accelerating simulations by spatially mixing interatomic potentials of different complexities. Through constrained linear fitting, we show it is possible to generate a ‘cheap’ approximate model which closely matches an ‘expensive’ reference in relevant regions of configuration space. We demonstrate the capability of ML-MIX through case-studies in Si, Fe, and W–He systems, achieving up to an 11× speedup on $\sim$ 8,000 atom systems without sacrificing accuracy on static and dynamic quantities, including calculation of minimum energy paths and dynamical simulations of defect diffusion. For larger domain sizes, we show that the achievable speedup of ML-MIX simulations is limited only by the relative speed of the cheap potential over the expensive potential. The ease of use and flexible nature of this method will extend the practical reach of MLIPs throughout computational materials science, enabling parsimonious application to large spatial and temporal domains.
@@ 1 Introduction
Atomistic simulations face two central challenges: accuracy in the model for atomic energies and forces and capturing physical processes at large length- and time-scales. Addressing each challenge under fixed computational budget leads to cost-accuracy trade-offs which have driven the development of many methods that approximate the underlying quantum-mechanical ground truth. The most important (and most successful) approximation for atomic interactions is density functional theory (DFT) — which pushed the size of simulations from one or two small atoms to hundreds or even thousands on modern machines — a meteoric development recognised with the 1998 Nobel Prize [1, 2].
@@
A further level of approximation is to do away with electrons completely; instead making predictions about atomic behavior purely from the positions of atomic nuclei and chemical species. For a many years, the most accurate models in this category were fit-by-hand interatomic potentials with simple functional forms designed to match experimental quantities of interest [3, 4, 5, 6] or DFT derived properties [7, 8]. Many such ‘empirical’ potentials still see considerable use today [9]. In the last three decades a new class of interatomic potentials has emerged, which have complex functional forms and parameters fit to DFT data, with an aim to reproduce large swathes of the potential energy surface [10, 11, 12, 13, 14, 15]. These machine learned interatomic potentials (MLIPs) have consistently achieved remarkable accuracy over a range of applications across materials science and chemistry [16, 17, 18, 19]. Whilst the evaluation costs of empirical potentials and MLIPs both scale linearly with the number of atoms in a system, the flexible functional forms of MLIPs result in much larger scaling pre-factors. Generally, MLIPs are approximately 2-3 orders of magnitude slower in evaluation than their simple empirical counterparts [20, 21].
@@
This cost has led to the uptake of MLIPs being slow in fields where simulation domains are prohibitively large and simulation timescales are long, for example investigations requiring molecular dynamics (MD) simulations of millions of atoms for billions of timesteps. These simulations can be split into two categories, those in which the entire domain is chemically complex, with atoms involved in bond-making/breaking processes throughout (e.g, amorphous materials [22, 23], liquids [24, 25], high entropy alloys [26, 27]), and those where the domain contains only isolated regions of complexity surrounded by atoms in simple environments (e.g, materials defects [28], biochemical simulations [29], catalysis [30]). In the second case, simulation cost can in principle be reduced by using cheap, approximate models away from the chemically ‘important’ regions of high complexity.
@@
This idea has been extensively explored in the context of quantum mechanical/ molecular mechanics modeling (QM/MM), a method which couples an expensive first-principles method (typically DFT) with a linear scaling potential for the complex (QM) and simple (MM) regions of a simulation domain respectively [31, 32, 33, 34, 35, 36, 37, 38, 39]. This paper details a method that is conceptually similar to QM/MM, but where the ‘QM’ and ‘MM’ regions are replaced with linear-scaling MLIPs of different complexities (ML/ML). The key difference is that ML/ML simulations aim to run for millions of timesteps, compared to only hundreds or thousands in QM/MM, with each timestep taking 3-6 orders of magnitude less wall time. ML/ML therefore incurs different software challenges: to minimise overhead, operations that take place each timestep must be embedded directly into the MD code. Whilst some studies have explored similar ‘ML/MM’ ideas [40, 41, 42], no generic, computationally parallel, simple to use implementation of ML/ML exists.
@@
In this paper we present ML-MIX, a package created to enable the whole workflow of accelerating a simulation within the popular open-source MD software LAMMPS [43, 44]. We show that if one starts with an ‘expensive’ MLIP, it is possible to generate a ‘cheap’ MLIP that locally approximates the potential energy surface, and this can be spatially mixed with the expensive MLIP to run accurate simulations at a fraction of the cost. The process is schematically depicted in Fig. 1.
The package is designed to be user-friendly; provided that a user has access to an expensive potential, our aim is that it should be straightforward to generate a cheap potential and attain a reasonable simulation speedup with minimal effort. To ensure the cheap potential locally matches the expensive potential, we generate the cheap potential through constrained linear fitting. The constraints we apply can be divided into two types depending on the strength of regularization they impose on the fit; ‘soft’ and ‘hard’. Soft constraint data covers the whole region of potential energy space that the cheap potential should approximate. Hard constraint data covers important properties which should match closely between potentials. The fit is constrained to a subspace in which predictions on the hard constraint data match near-exactly between the cheap and expensive potentials. The constrained fitting procedure is described in detail in section 4.1.
[…]
@@
Figure 1: A schematic overview of the simulation acceleration workflow. Left hand side: the constrained potential fitting process, whereby an accurate expensive potential (orange) is approximated in local regions of potential energy space by a simpler cheap potential (blue). Right hand side: schematic showing how simulation is accelerated; cheap potential is used to evaluate force components on atoms in ‘bulk-like’ environments.
@@ 2.1 Potential fitting
The starting point for our study are three ‘expensive’ linear atomic cluster expansion (ACE) potentials, each of which was fit to expansive and publically available DFT datasets (Si [16], Fe [45], W–He [46]) using the linear ACEpotentials.jl package [47]. The Fe and W–He potentials were of correlation order 3 and maximum total polynomial degree 20, whilst the Si potential had a larger basis set, with correlation order 4 and maximum total polynomial degree 20. The Si potential cutoff was 6.0 Å, the Fe potential cutoff was 5.5 Å, and the W–He potential cutoff was 5.0 Å. For the Si potential, the potential parameters and dataset match those described in [47]. For the Fe and W–He potentials, full sets of fitting parameters are given in appendix A.
@@
These expensive potentials were used to generate ‘synthetic’ training data focused around a suitable local region of potential energy space. In each case, soft constraint data was a bulk-crystal high temperature MD trajectory, and hard constraint data was a set of zero-temperature homogeneous lattice deformations. Energy and force evaluations were performed using the ML-PACE package [20]. As the energy of weak, homogeneously deformed lattice structures are described fully by the elastic constants [38], the selected hard constraint enforced elastic constant matching between potentials. Additional details on the fitting of these potentials is presented in section 4.1.2.
@@
Table 1: RMSE table for both expensive and cheap potentials on the full DFT datasets used for the creation of the expensive potentials and the small synthetic MD datasets used for the creation of the cheap potentials. For the W-He potential, only configurations containing W are used (as the cheap UF3 potential cannot describe He interactions). Note that the synthetic data is generated by the expensive potential, hence the lack of RMSE values in the lower left quadrant.
Expensive potentials Cheap potentials
@@
Full DFT data Si 2.14 0.075 2022.68 0.79
Fe 2.32 0.083 152.15 0.30
W 2.15 0.070 11938.00 157.54
Synthetic MD data Si - - 0.21 0.045
Fe - - 0.81 0.090
W - - 1.02 0.053
Cheap potential models: For Si and Fe, the cheap potentials were small ACE potentials, with correlation orders of 2 and maximum total polynomial degrees of 10, with potential cutoffs matching those of the reference models. For the W-He simulation, the cheap potential was a pure W ultra-fast (UF3) potential [21] with a two-body cutoff of 6.0 Å and a 3 body cutoff of 3.5 Å. The energy and force root mean squared error (RMSE) values for all potentials evaluated on both the expensive potential DFT datasets and cheap potential synthetic datasets are displayed in Table. 1. The cheap potentials perform very badly compared to the expensive potential when tested across the whole potential energy surface, but exhibit excellent performance when evaluated in their target domain.
A plot for the relative speedups of each cheap potential over the expensive potentials on different numbers of processors are shown in Fig. 2. Data is shown for a small domain ($\sim$8k atoms) and a large domain ($\sim$250k atoms), with the speedup curves being approximately the same for both. The serial speedup for the Si potential is $\sim$19x, but this increases up to $\sim$30x in parallel. Similar behaviour is seen for the W–He potentials, the relative speedup is $\sim$59x in serial, but in parallel increases up to $\sim$75x. In contrast, the Fe cheap potential remains $\sim$8x faster in both serial and parallel. It is unclear exactly why the W–He and Si cheap potentials show this additional parallel speedup, and is likely due to the implementation of the ML-PACE package [20].
Figure 2: Speedup of simulations using only the expensive potential relative to the same simulations using only the cheap potential on 8k and 250k atoms with an increasing number of MPI parallel domains. Trendlines are given by a Gaussian process fit to all the datapoints in each group. It can be seen that the W–He and Si cheap potentials exhibit a greater speedup over the expensive potentials in parallel than serial. Error bars on each data point represent the standard deviation in the measured speedups over three repeats.
@@
Table 2 shows the impact of the hard constraint on the elastic constants and lattice parameters predicted by the cheap potentials. The elastic constants were computed using the matscipy package [48]. The unconstrained fit (fitting to MD data only) produced potentials with mismatched elastic constants and lattice parameters relative to the expensive potential reference. In the case of the Si and Fe ACE potentials, adding the hard constraint data led to near perfect matching of the reference elastic constants. For the pure W UF3 potential the fitting was less successful; it was determined the UF3 functional form was insufficiently flexible to match the hard constraint data and simultaneously give a low RMSE on the MD data. Under the constraint, the $\mathrm{C}_{11}$ and $\mathrm{C}_{44}$ elastic constants got closer to the underlying reference, but $\mathrm{C}_{12}$ became worse. In all cases, the equilibrium lattice parameter of the model became more accurate under constrained fitting. The RMSE in the predicted energies and forces on the MD data for both the constrained and unconstrained fits are shown in table 3. Forcing the fit to lie in the constraint subspace always increases the RMSE.
@@
Table 2: Table of elastic constants and lattice parameters of according to each potential (expensive, cheap unconstrained, cheap constrained) for Si, Fe and W–He. Elastic constants are fit to stress-strain data using Bayesian linear regression, with the error $\sigma$ in each value found from the maximum likelihood estimate. For Si and Fe, the addition of the hard constraints to the cheap potential fitting produces elastic constants which match the expensive within $1\sigma$. For the W–He potential, the constrained fitting improves $C_{11}$ (matching within $1\sigma$) and $C_{44}$ (matching within $2\sigma$), but leads to a significantly worse $C_{12}$ than the unconstrained case. In all cases, the constraint dramatically improves the lattice parameter.
$C_{11}$ (GPa) $C_{12}$ (GPa) $C_{44}$ (GPa) a (Å)
Si Expensive (reference) 150.8 $\pm$ 1.0 56 $\pm$ 2 70.5 $\pm$ 0.4 5.461
Cheap (unconstrained) 139 $\pm$ 3 81.29 $\pm$ 0.12 52.2 $\pm$ 1.9 5.661
Cheap (constrained) 148 $\pm$ 2 55 $\pm$ 3 69.8 $\pm$ 0.3 5.461
Fe Expensive (reference) 283 $\pm$ 12 154 $\pm$ 6 105 $\pm$ 6 2.834
Cheap (unconstrained) 426 $\pm$ 16 364 $\pm$ 2 239.0 $\pm$ 1.9 3.360
Cheap (constrained) 284 $\pm$ 11 154 $\pm$ 7 106 $\pm$ 3 2.834
W—He Expensive (reference) 520 $\pm$ 14 199 $\pm$ 5 143 $\pm$ 5 3.181
Cheap (unconstrained) 604 $\pm$ 12 209 $\pm$ 12 180 $\pm$ 6 3.136
Cheap (constrained) 527 $\pm$ 11 157 $\pm$ 4 157 $\pm$ 3 3.180
Table 3: RMSE table for the constrained and unconstrained cheap potentials on the synthetic MD dataset. The RMSEs get slightly worse when the fit is constrained.
Unconstrained fit Constrained fit
Energy
Force
(eV/Å)
Energy
Force
(eV/Å)
Si 0.15 0.038 0.21 0.045
Fe 0.14 0.083 0.81 0.090
W–He 0.12 0.053 1.02 0.053

### v2

@@ Abstract
Machine-learned interatomic potentials can offer near first-principles accuracy but are computationally expensive, limiting their application to large-scale molecular dynamics simulations. Inspired by quantum mechanics/molecular mechanics methods we present ML-MIX, a CPU- and GPU-compatible LAMMPS package to accelerate simulations by spatially mixing interatomic potentials of different complexities allowing deployment of modern MLIPs even under restricted computational budgets. We demonstrate our method for ACE, UF3, SNAP and MACE potential architectures and demonstrate how linear ‘cheap’ potentials can be distilled from a given ‘expensive’ potential, allowing close matching in relevant regions of configuration space. The functionality of ML-MIX is demonstrated through tests on point defects in Si, Fe and W–He, in which speedups of up to 11× over $\sim$ 8,000 atoms are demonstrated, without sacrificing accuracy. The scientific potential of ML-MIX is demonstrated via two case studies in W, measuring the mobility of $\it{b}$ = $\frac{1}{2}\langle 111\rangle$ screw dislocations with ACE/ACE mixing and the implantation of He with MACE/SNAP mixing. The latter returns He reflection coefficients which (for the first time) match experimental observations up to an He incident energy of 80 eV – demonstrating the benefits of deploying state-of-the-art models on large, realistic systems.
@@ 1 Introduction
Atomistic simulations face two central challenges: modeling quantum-mechanical atomic energies and forces and capturing physical processes at large length- and time-scales. Addressing these challenges under fixed computational budget leads to cost-accuracy trade-offs which have driven methodological development over many decades. The most important (and most successful) approximation for electronic interactions is density functional theory (DFT) – a mean field approach which increases the size of simulations from one or two small atoms to hundreds or even thousands on modern machines [Burke_2012, Jones_2015].
@@
A further level of approximation is to treat electrons implicitly, predicting atomic energies and forces only from nuclear positions and chemical species. For many years, the most accurate models in this category were interatomic potentials with simple functional forms fit to match experimental quantities of interest [Daw_1984, Lennard-Jones_1931, Stillinger_1985, Tersoff_1988] or DFT derived properties [Ercolessi_1994, Brommer_2007]. Many such ‘empirical’ potentials still see considerable use today [Muser_2023]. In the last three decades a new class of interatomic potentials has emerged, which have complex functional forms and parameters fit to DFT data, with an aim to reproduce large swathes of the potential energy surface [Behler_2007, Bartók_2010, Shapeev_2016, Drautz_2019, Deringer_2021, Batatia_2022]. These machine learned interatomic potentials (MLIPs) have achieved remarkable accuracy over a range of applications across materials science and chemistry [Bartók_2018, Kovács_2023, Qamar_2023, Batatia_2024]. Whilst the evaluation costs of empirical potentials and MLIPs both scale linearly with the number of atoms in the system, the complex functional forms of MLIPs result in much larger scaling pre-factors; MLIPs are typically 2-3 orders of magnitude slower in evaluation time than their simple empirical counterparts [Lysogorskiy_2021, Xie_2023].
@@
This cost has slowed the uptake of MLIPs in fields where simulation domains are large and/or timescales are long, which encompasses a large range of phenomena in materials science. It is useful to split such simulations into two categories, those where the entire domain is chemically complex, with atoms involved in bond-making/breaking processes throughout (e.g, amorphous materials [Qian_2023, Grießer_2023], liquids [Al-Awad_2023, Tian_2023], high entropy alloys [Byggmästar_2021, Shiotani_2023]), and those where the domain contains only isolated regions of complexity surrounded by atoms in simple environments (e.g, materials defects [Gleizer_2014], biochemical simulations [Cui_2021], catalysis [Bramley_2023]). In the second case, simulation cost can in principle be reduced through spatial decomposition: using cheaper, approximate models away from the chemically ‘important’ regions of high complexity.
@@
Spatial decomposition has been extensively explored in the context of quantum mechanical/ molecular mechanics modeling (QM/MM), a method which couples an expensive first-principles method (typically DFT) with a linear scaling potential for the complex (QM) and simple (MM) regions of a simulation domain respectively [Warshel_1976, Chen_Ortner_2016, Chen_Ortner_2017, Swinburne_and_Kermode_2017, Bernstein_2009, Kermode_2008, Molani_2024, Grigorev_2023, Wagih_2022]. In this paper, we detail a method that is conceptually similar to QM/MM, but instead combines MLIPs of different complexities (ML/ML). The key difference is that ML/ML simulations aim to run for millions of timesteps, compared to only hundreds or thousands in QM/MM, with each timestep taking 3-6 orders of magnitude less wall time. ML/ML therefore incurs different software challenges: to minimise overhead, operations that take place each timestep must be closely integrated into the molecular dynamics (MD) software. Whilst some recent studies have explored similar ‘ML/MM’ ideas [Hofstetter_2022, Böselt_2021, Immel_2024, Galvelis_2023], no generic implementation of ML/ML exists.
@@
The central contribution of this work is ML-MIX, a package created to enable the whole workflow of accelerating a simulation within the popular open-source MD software LAMMPS [LAMMPS, type_label_LAMMPS]. Given an ‘expensive’ MLIP, we demonstrate how one can distill a ‘cheap’ linear MLIP that locally approximates the potential energy surface and then show how spatial mixing of the cheap and expensive MLIPs can be used to run accurate simulations at a fraction of the cost. The process is schematically depicted in Fig. 1. ML-MIX aims to be as generic as possible; it is able to perform ML/ML and ML/MM simulations with many native LAMMPS potentials out of the box, in parallel, on both CPU and GPU. A non-exhaustive list of compatible MLIPs includes SNAP [Thompson_2015], ACE [Drautz_2019], UF3 [Xie_2023] and MACE [Batatia_2022]. An up-to-date full list of tested compatible potentials can be found in the ML-MIX GitHub repository [ML-MIX].
Here we present the results of six ML/ML case-studies, targeting one static and five dynamic quantities of interest in Si, Fe and W, demonstrating that ML-MIX can be used to generate results that agree with those from fully expensive simulations (and in the final case-study, experiment) at a fraction of the cost. The first case studies are designed primarily to demonstrate different features of the ML-MIX package. These particularly aim to highlight how ML/ML methods can produce accurate results for simple quantities, even for very small ($\sim 50$ basis function) linear cheap potentials. We then present two more case studies which were primarily conducted for scientific interest. Both studies concern the application of W as a diverter material in nuclear fusion reactors. In the first of these final two case-studies, we consider the thermally activated glide of screw dislocations in W at different stresses and temperatures. In the second we consider the normally incident deposition of He at different energies into W at 1000 K. In this final case-study, we were able to reproduce (for the first time) experimental observations of the reflection coefficient of He into W for deposition energies up to 80 eV.
@@
Figure 1: A schematic overview of the simulation acceleration workflow. Left side: the constrained potential fitting process, whereby an accurate expensive potential (orange) is approximated in local regions of potential energy space by a simpler cheap potential (blue). Right side: schematic showing how simulation is accelerated; cheap potential is used to evaluate force components on atoms in ‘bulk-like’ environments.
@@ 2.1 Potential fitting
The starting point for our study are five ‘expensive’ potentials; four linear atomic cluster expansion (ACE) potentials (Si, Fe, W, W – He) fit using the ACEpotentials.jl package [Witt_2023], and one MACE potential created from fine-tuning the mpa-0-medium foundation model to W – He data using multihead replay fine-tuning [Batatia_2024]. Table 1 shows the fitting parameters of each expensive potential and where to find the corresponding DFT data and parameters. Additional hyper-parameters for the Fe, W and W – He ACE potentials are specified in section S1 of the supplementary information. The root-mean squared per-atom-energy and force errors (RMSEs) for all potentials on their fitting datasets are shown in the upper left quadrant of Table 2.
Table 1: A summary of expensive potential fitting parameters and source of corresponding DFT data. For the ACE potentials, $\nu$ refers to correlation order and $p$ refers to maximum total polynomial degree. For the MACE potential, $L$ refers to the message rank. In all cases, $r_{\mathrm{cut}}$ is the cutoff radius for the local environment.
Expensive
potential
Key fitting
parameters
DFT
Si ACE
$\nu=4$
$p=20$
$r_{\mathrm{cut}}=6.0$ Å
[Bartók_2018]
Fe ACE
$\nu=3$
$p=20$
$r_{\mathrm{cut}}=5.5$ Å
[Zhang_2023]
W–He ACE
$\nu=3$
$p=20$
$r_{\mathrm{cut}}=5.0$ Å
[Nutter_2024, Nutter_2024_supp]
W ACE
$\nu=3$
$p=21$
$r_{\mathrm{cut}}=5.0$ Å
Data: [ml-mix-zenodo]
Params: [Nutter_2024, Nutter_2024_supp]
W–He MACE
mpa-0-medium [Batatia_2024]
fine tuned
128 L=1 channels
128 L=0 channels
2 layers
$r_{\mathrm{cut}}=6.0$ Å
Pretraining data:
MPtraj [mptraj]
sAlex [salex_cite_1, salex_cite_2]
Fine tuning:
Data: [Nutter_2024, Nutter_2024_supp, ml-mix-zenodo]
Params: [Nutter_2024, Nutter_2024_supp]
The following cheap potentials were generated: (i-ii) Si and Fe linear ACE potentials (fitted with ACEpotentials.jl) with correlation orders of 2 and maximum total polynomial degrees of 10 and cutoffs of 6.0 Å and 5.5 Å respectively, (iii) a W ultra-fast (UF3 [Xie_2023]) potential with a two-body cutoff of 6.0 Å and a 3-body cutoff of 3.5 Å and (iv) A W linear ACE potential (fitted with ACEpotentials.jl) with a correlation order of 3 and a maximum polynomial degree of 14 and a cutoff of 5.0 Å. We also use (v) the pure W SNAP potential generated by Wood and Thompson [wood_2017]. Whilst all these potentials have significantly lower evaluation cost than their expensive counterparts, (iv) and (v) are more sophisticated than (i-iii). This is due to (i-iii) only needing to describe relatively simple bulk interactions, whilst (iv) and (v) need to also describe surfaces and longer-range strain fields. In each case, cutoff radii were selected for the cheap potential which were the same or smaller than the expensive potential.
The correlation order 3, maximum total polynomial degree 14 W ACE was fit to the same DFT dataset as its expensive counterpart (the correlation order 3 maximum polynomial degree 21 W ACE). RMSE values for this fit are shown on upper right quadrant of Table 2. Whilst the W SNAP potential was also fit to much of the same underlying DFT data [wood_2017], the full fitting dataset is unpublished and therefore the RMSEs are left out of Table 2.
For $\sim 5000$ atoms, the speedup for the 3, 14 W ACE over the 3, 21 W ACE was $\sim 4.4\times$, whilst the speedup of the pure W SNAP potential over the fine-tuned mpa-0-medium model was $\sim 29\times$. Note that the MACE/SNAP speedup was measured on GPU, whilst all other quoted speedups were measured on CPU.
@@
To generate the small cheap potentials (Si and Fe ACE, W UF3), the expensive potentials were used to generate ‘synthetic’ training data focused around a suitable local region of the potential energy surface. In each case, the synthetic fitting data was divided into two types, ‘soft constraint’ data which consisted of bulk-crystal high-temperature MD trajectories, and ‘hard constraint’ data which comprised a set of zero-temperature homogeneous lattice deformations. Constrained linear fitting was used to ensure that elastic constants and lattice parameters closely matched between the cheap and expensive potentials. The constrained fitting scheme is described in detail in section 4.1.
@@
Table 2: Training RMSE table for both expensive and cheap potentials on the full DFT datasets used for the creation of the expensive potentials and the small synthetic MD datasets used for the creation of the cheap potentials. For the W-He potential, only configurations containing W are used (as the cheap UF3 potential cannot describe He interactions). Note that the synthetic data is generated by the expensive potential, hence the lack of RMSE values in the lower left quadrant. Energy errors are per-atom. The full DFT data used for the fitting of the the W SNAP is not public, so these RMSEs are not stated.
RMSEs RMSEs
Expensive
potentials
@@
Full DFT data Si ACE (4, 20) 2.14 0.075 Si ACE (2, 10) 106.5 0.27
Fe ACE (3, 20) 2.32 0.083 Fe ACE (2, 10) 18.48 0.156
W - He ACE (3, 20) 2.15 0.070 W UF3 6549 0.16
W ACE (3, 21) 1.90 0.066 W ACE (3, 14) 4.489 0.116
W - He MACE 4.10 0.075 W SNAP * *
Synthetic MD data Si ACE (4, 20) - - Si ACE (2, 10) 0.21 0.045
Fe ACE (3, 20) - - Fe ACE (2, 10) 0.81 0.090
W - He ACE (3, 21) - - W UF3 1.02 0.053
The training energy and force root mean squared error (RMSE) values for the synthetically trained potentials are displayed in the lower right quadrant of Table. 2. For comparison, potentials with identical hyper-parameters were also fit directly to the same DFT data as the expensive reference. The training RMSE for these fits are shown in the upper right quadrant of Table. 2. Compared to the corresponding expensive potentials, the speedup for force evaluations with the Si ACE potential was $\sim 30\times$, the W UF3 potential was $\sim 75\times$, and the Fe ACE potential was $\sim 8\times$. Measured speedups were the same for systems of both 8000 and 250,000 atoms.
@@
Table 3 shows the elastic constants predicted by the expensive and cheap potentials. For the potentials fit to synthetic data, it also includes rows demonstrating the impact of applying the hard constraint on the elastic constants and lattice parameters. The elastic constants were computed using the matscipy package [matscipy]. In the case of the Si and Fe ACE potentials, adding the hard constraint data led to near-perfect matching of the reference elastic constants. For the pure W UF3 potential, the $\mathrm{C}_{11}$ and $\mathrm{C}_{44}$ elastic constants got closer to the underlying reference, but $\mathrm{C}_{12}$ was worse, something that can be attributed to the simplicity of the UF3 functional form. In all cases, the equilibrium lattice parameter of the model became closer to the expensive reference under constrained fitting.
@@
Table 3: Elastic constants and lattice parameters according to each potential (expensive, cheap unconstrained, cheap constrained) for Si, Fe and W–He. Elastic constants are fit to stress-strain data using Bayesian linear regression, with the error ($\sigma$) in each value found from the maximum likelihood estimate. For Si and Fe, the addition of the hard constraints to the cheap potential fitting produces elastic constants which match the expensive within $1\sigma$. For the W–He potential, the constrained fitting improves $\mathrm{C}_{11}$ (matching within $1\sigma$) and $\mathrm{C}_{44}$ (matching within $2\sigma$), but leads to a significantly worse $\mathrm{C}_{12}$ than the unconstrained case. In all cases, the constraint dramatically improves the lattice parameter.
$\mathrm{C}_{11}$ (GPa) $\mathrm{C}_{12}$ (GPa) $\mathrm{C}_{44}$ (GPa) a (Å)
Si ACE (4, 20) Expensive 150.8 $\pm$ 1.0 56 $\pm$ 2 70.5 $\pm$ 0.4 5.461
Si ACE (2, 10) Cheap (unconstrained) 139 $\pm$ 3 81.29 $\pm$ 0.12 52.2 $\pm$ 1.9 5.661
Si ACE (2, 10) Cheap (constrained) 148 $\pm$ 2 55 $\pm$ 3 69.8 $\pm$ 0.3 5.461
Fe ACE (3, 20) Expensive 283 $\pm$ 12 154 $\pm$ 6 105 $\pm$ 6 2.834
Fe ACE (2, 10) Cheap (unconstrained) 426 $\pm$ 16 364 $\pm$ 2 239 $\pm$ 2 3.360
Fe ACE (2, 10) Cheap (constrained) 284 $\pm$ 11 154 $\pm$ 7 106 $\pm$ 3 2.834
W–He ACE (3, 20) Expensive 520 $\pm$ 14 199 $\pm$ 5 143 $\pm$ 5 3.181
W UF3 Cheap (unconstrained) 604 $\pm$ 12 209 $\pm$ 12 180 $\pm$ 6 3.136
W UF3 Cheap (constrained) 527 $\pm$ 11 157 $\pm$ 4 157 $\pm$ 3 3.180
W ACE (3, 21) Expensive 520 $\pm$ 14 199 $\pm$ 5 143 $\pm$ 5 3.180
W ACE (3, 14) Cheap (unconstrained) 520 $\pm$ 14 196 $\pm$ 5 144 $\pm$ 5 3.180
W–He MACE
Expensive 540 $\pm$ 15 215 $\pm$ 8 147 $\pm$ 4 3.179
W SNAP Cheap (unconstrained) 518 $\pm$ 14 196 $\pm$ 5 144 $\pm$ 5 3.180